# Rethinking Optimal Verification Granularity for Compute-Efficient Test-Time Scaling

### v1

@@ Abstract
Test-time scaling (TTS) has proven effective in enhancing the reasoning capabilities of large language models (LLMs). Verification plays a key role in TTS, simultaneously influencing (1) reasoning performance and (2) compute efficiency, due to the quality and computational cost of verification. In this work, we challenge the conventional paradigms of verification, and make the first attempt toward systematically investigating the impact of verification granularity—that is, how frequently the verifier is invoked during generation, beyond verifying only the final output or individual generation steps. To this end, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes beam search and Best-of-N sampling via a tunable granularity parameter $g$. Extensive experiments with VG-Search under varying compute budgets, generator-verifier configurations, and task attributes reveal that dynamically selecting $g$ can improve the compute efficiency and scaling behavior. Building on these findings, we propose adaptive VG-Search strategies that achieve accuracy gains of up to 3.1% over Beam Search and 3.6% over Best-of-N, while reducing FLOPs by over 52%. We will open-source the code to support future research.
@@ 1 Introduction
Figure 1: Motivations for Optimizing Verification Granularity $g$. (a) Small PRM (Llama3.1-8B-PRM) score deltas across generation steps indicate redundant verification. (b) Verification incurs notable latency overheads. (c) Increasing verification granularity lowers the latency of both Proposer and Verifier. (d) Optimizing $g$ improves the accuracy-compute tradeoff over fixed-granularity baselines (Qwen2.5-Math-7B as generator, Skywork-o1-1.5B as verifier).
@@
To explore RQ1, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes verifier-guided Beam Search and Best-of-N methods. VG-Search employs a granularity parameter $g$ that controls the verification granularity (Section 2). Our hardware performance profiling (Figure 1(c)) shows that $g$ has a significant impact on the latency contributions of both the generator LLM and the verifier. To further investigate the accuracy–compute trade-off under varying $g$, we conduct extensive experiments with VG-Search across different values of $g$, models, compute budget, and datasets (Section 4). Building on top of the insights from RQ1, we propose an adaptive VG-Search approach which dynamically adjusts the verification granularity based on the compute budget and task attributes (Section 5) to further address RQ2. As shown in Figure 1(d), our results demonstrate better performance-compute frontiers than previous sampling-based TTS methods. Overall, our main contributions are:
@@ 2.1 Verification Granularity in Verifier-Guided Search
The goal of the search process is to generate a correct or high-quality solution $\mathbf{s}_{1:T}^{*}$ of length $T$ via interaction between the generator and verifier. A key hyperparameter in this process is the verification granularity $g$, which controls how frequently the verifier can intervene and guide the generator. To illustrate the impact of $g$, we link the verifier-guided search to the Infinite Monkey Theorem (IMT) [38], where a random generator (monkey) tries to reproduce a target sequence (e.g., Shakespeare’s work) under a perfect verifier invoked every $g$ characters. When verification is infrequent (large $g$), the expected number of attempts grows exponentially as $\mathbb{E}[A^{g}]$, making success computationally infeasible. In contrast, with frequent verification, early error pruning reduces the expected cost $\mathbb{E}[A^{g}]$, at the expense of higher verification overhead. This toy setting underscores a fundamental trade-off between verification cost and exploration efficiency. Optimizing $g$ is thus critical for effective collaboration between generator $\mathcal{G}$ and verifier $\mathcal{V}$ under compute constraints.
@@ 2.2 Variable-Granularity Search
    Extend: For each of the $B_{1}$ selected beams, produce $g-1$ generation steps using $\mathcal{G}$.
@@
Omitting Step 3 (Extend) or setting $g=1$ reduces VG-Search to standard Beam Search. At the other extreme, when $g=L$ (the number of generation steps in a full solution), VG-Search becomes equivalent to Best-of-N sampling. One alternative design is to produce $g-1$ generation steps in Branch, mimicking Beam Search but with variable verification granularity. However, prior work [32] shows that rolling out multiple future tokens, as in LookAhead Search, does not yield the best performance. We therefore adopt a simpler, more compute-efficient design: Branch generates just one step per candidate, and early pruning via the verifier occurs before Extend. As a result, only the top $B_{1}$ candidates continue in Extend, instead of the full $B_{1}\times B_{2}$ set used in Beam Search or Best-of-N—leading to significantly lower proposer FLOPs. Larger $g$ further amplifies savings for both verifier and proposer. We validate this design in Section 4.1. A case study of VG-Search is illustrated in Appendix A.5.
@@ 4.2 Trade-offs: Granularity, Verifier Parameter, and Branching
To assess the practical limits of sparse verification, we define the largest effective $g$ as the maximum verification granularity that retains at least 95% of the accuracy achieved at $g{=}1$. We also analyze the accuracy gain of $g{=}2$ over $g{=}1$ across varying sample budgets and task difficulty levels. Figure 4(d),  4(e) show that the optimal $g$ is strongly influenced by both difficulty and sampling budget.
Increased Task Difficulty Demands Denser Verification. Across all models and compute budgets, we observe a consistent trend: as task difficulty increases, the largest effective $g$ decreases. For the hardest problems (Levels 4–5), $g{=}1$ or $g{=}2$ is often required to maintain performance at lower compute budget. Likewise, the advantage of $g{=}2$ over $g{=}1$ diminishes with difficulty, likely because harder tasks involve more reasoning errors, necessitating more frequent verifier intervention to stay on track.
Higher Number of Samples Tolerates Sparser Verification. In general, larger sample budgets allow for sparser verification, especially on easy to moderately difficult tasks. Higher budgets also amplify the performance gain of $g{=}2$ over $g{=}1$, as the generator can explore a wider search space, increasing the likelihood of finding a correct path despite less frequent verification.
@@ 5 Towards Adaptive Redefinition of the Verification Granularity
Our results (Figures 3, 4) show that the optimal $g$ in VG-Search depends on task difficulty, compute budget, and model capability. This motivates a more systematic approach to defining the search step, addressing RQ2. We formulate an optimization dual for selecting verification granularity: minimizing compute for a target accuracy vs. maximizing accuracy under a compute budget. For each, we propose a corresponding adaptive granularity strategy:
@@
We compare against Beam Search, DVTS and Best-of-N baselines on the MATH-500 test set. "val" indicates $g$ was tuned on a validation set MATH-250, while strategies under "oracle" use oracle $g$ selected on the test set. We set $\epsilon=0$ for CM-$g$. Table 1 shows that both AM-$g$ and CM-$g$ improve performance and efficiency. AM-$g$ consistently achieves higher accuracy, with gains up to 3.1% over Beam Search and 3.6% over Best-of-N, while CM-$g$ provides substantial FLOPs savings, reducing compute by over 50% at $n=64$ and $n=256$ while maintaining or improving accuracy. For instance, CM-$g$ (val) at $n=128$ achieves 89.9% accuracy using just 11086 FLOPs—only 46% of the baseline budget. Although oracle-tuned variants perform slightly better, validation-tuned versions (CM-$g$ (val), AM-$g$ (val)) still outperform fixed-$g$ methods, demonstrating strong generalization and practicality. Under our cost model, $C_{\mathcal{G}}$ dominates, so most FLOPs savings come from pruning candidate paths during the Extend Step (Section 2.2). In summary, adapting verification granularity $g$ to task difficulty and compute budget offers a simple yet effective way, enabling more efficient and performant LLM reasoning.
@@ A.2 FLOPs Calculation using the Cost Model
where $L$ is average solution length in generation steps, $T$ is the average number of tokens per generation step. To simplify compute tracking across experiments, we define a normalized compute proxy:
@@ A.3 Experiment Details for VG-Search
Baselines and Voting Methods. Our codebase is based on HuggingFace’s search-and-learn [4]. We adopt three mainstream TTS strategies: Best-of-N [5], Beam Search [32] and DVTS [4]. For all strategies we explore number of samples of 4, 16, 64, 128 and 256. For Beam Search and DVTS we use a fixed Branch-Out Factor $B_{2}=4$ following [32]. For all strategies, we use majority voting [37]. Temperature is set to 0.8 for all problems following HuggingFace’s setup.
@@
Hardware and Frameworks. We conduct our experiments on NVIDIA H100 [12] and A100 [11] GPUs, using CUDA 12.8 on Ubuntu 22.04. The vLLM library (v0.6.3) [18] is employed for model execution. The experiments for VG-Search on MATH-500 take from 1 to 5 hours, depending on the candidate number. To obtain accurate system-level runtime measurements, we use Nsight Systems [26], and employ the NVIDIA NVTX [13] extension to categorize execution time across different models. For the latency measurement in Figure 1(c), we use Qwen2.5-Math-1.5B as generator, while using RLHFlow/Llama3.1-8B-PRM-Deepseek-Data as the verifier.
A.4 Limitations
@@
A.5 Case Study of VG-Search

### v2

@@ Abstract
Test-time scaling (TTS) has proven effective in enhancing the reasoning capabilities of large language models (LLMs). Verification plays a key role in TTS, simultaneously influencing (1) reasoning performance and (2) compute efficiency, due to the quality and computational cost of verification. In this work, we challenge the conventional paradigms of verification, and make the first attempt toward systematically investigating the impact of verification granularity—that is, how frequently the verifier is invoked during generation, beyond verifying only the final output or individual generation steps. To this end, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes beam search and Best-of-N sampling via a tunable granularity parameter $g$. Extensive experiments with VG-Search under varying compute budgets, generator-verifier configurations, and task attributes reveal that dynamically selecting $g$ can improve the compute efficiency and scaling behavior. Building on these findings, we propose adaptive VG-Search strategies that achieve accuracy gains of up to 3.1% over Beam Search and 3.6% over Best-of-N, while reducing FLOPs by over 52%. Our code is avaiblae at github.com/hmarkc/VG-Search.
@@ 1 Introduction
Figure 1: Motivations for Optimizing Verification Granularity $g$. (a) Small PRM score deltas (LLaMA3.1-8B-PRM) across generation steps indicate redundant verification. (b) Verification incurs notable latency overheads. (c) Increasing verification granularity lowers the latency of both Proposer and Verifier. (d) Optimizing $g$ improves the accuracy-compute tradeoff over fixed-granularity baselines (Qwen2.5-Math-7B as generator, Skywork-o1-1.5B as verifier).
@@
To explore RQ1, we introduce Variable Granularity Search (VG-Search), a unified algorithm that generalizes verifier-guided Beam Search and Best-of-N methods. VG-Search employs a granularity parameter $g$ that controls the verification granularity (Section 2). Our hardware performance profiling (Figure 1(c)) shows that $g$ has a significant impact on the latency contributions of both the generator LLM and the verifier. To further investigate the accuracy–compute trade-off under varying $g$, we conduct extensive experiments with VG-Search across different values of $g$, models, compute budget, and datasets (Section 4). Building on top of the insights from RQ1, we propose an adaptive VG-Search approach which dynamically adjusts the verification granularity based on the compute budget and task attributes (Section 5) to further address RQ2. As shown in Figure 1(d), our results demonstrate better performance–compute frontiers than previous sampling-based TTS methods. Overall, our main contributions are:
@@ 2.1 Verification Granularity in Verifier-Guided Search
The goal of the search process is to generate a correct or high-quality solution $\mathbf{s}_{1:T}^{*}$ of length $T$ via interaction between the generator and verifier. A key hyperparameter in this process is the verification granularity $g$, which controls how frequently the verifier can intervene and guide the generator. To illustrate the impact of $g$, we link the verifier-guided search to an edge case, the Infinite Monkey Theorem (IMT) [38], where a random generator (monkey) tries to reproduce a target sequence (e.g., Shakespeare’s work) under a perfect verifier invoked every $g$ characters. When verification is infrequent (large $g$), the expected number of attempts grows exponentially as $\mathbb{E}[A^{g}]$, making success computationally infeasible. In contrast, with frequent verification, early error pruning reduces the expected cost $\mathbb{E}[A^{g}]$, at the expense of higher verification overhead. This toy setting underscores a fundamental trade-off between verification cost and exploration efficiency. Optimizing $g$ is thus critical for effective collaboration between generator $\mathcal{G}$ and verifier $\mathcal{V}$ under compute constraints.
@@ 2.2 Variable-Granularity Search
    Extend: For each of the $B_{1}$ selected beams, produce g1g-1 generation steps using $\mathcal{G}$.
@@
Omitting Step 3 (Extend) or setting $g=1$ reduces VG-Search to standard Beam Search. At the other extreme, when $g=L$ (the number of generation steps in a full solution), VG-Search becomes equivalent to Best-of-N sampling. One alternative design is to produce g1g-1 generation steps in Branch, mimicking Beam Search but with variable verification granularity. However, prior work [32] shows that rolling out multiple future tokens, as in LookAhead Search, does not yield the best performance. We therefore adopt a simpler, more compute-efficient design: Branch generates just one step per candidate, and early pruning via the verifier occurs before Extend. As a result, only the top $B_{1}$ candidates continue in Extend, instead of the full $B_{1}\times B_{2}$ set used in Beam Search or Best-of-N—leading to significantly lower proposer FLOPs. Larger $g$ further amplifies savings for both verifier and proposer. We validate this design in Section 4.1. A case study of VG-Search is illustrated in Appendix A.8.
@@ 4.1 Test-Time Scaling Law with Verification Granularity
4.2 Ablation Studies on Verifier Quality and Model Scale
We conducted two ablation studies to investigate the impact of verifier quality and model scale on the optimal granularity, with results shown in Figure 4. To simulate a worst-case scenario for verifier quality, we replaced the standard verifier with a maximally noisy PRM that assigned a random score to each generation step. As hypothesized, when the verifier provides no useful signal, the optimal strategy consistently shifts to sparser verification ($g>1$), as frequent but unreliable guidance degrades search performance (Figure 4).
We validated our findings at different model scales by pairing various generators from the Qwen-2.5 series with Skywork-PRM verifiers of different sizes. The results (Figure 4) confirm that the optimal verification granularity remains dynamic across all tested scales.
Figure 4: Ablation on verifier quality and model sizes.
@@ 4.3 Trade-offs: Granularity, Verifier Parameter, and Branching
To assess the practical limits of sparse verification, we define the largest effective $g$ as the maximum verification granularity that retains at least 95% of the accuracy achieved at g=1g{=}1. We also analyze the accuracy gain of g=2g{=}2 over g=1g{=}1 across varying sample budgets and task difficulty levels. Figure 5(d),  5(e) show that the optimal $g$ is strongly influenced by both difficulty and sampling budget.
Increased Task Difficulty Demands Denser Verification. Across all models and compute budgets, we observe a consistent trend: as task difficulty increases, the largest effective $g$ decreases. For the hardest problems (Levels 4–5), g=1g{=}1 or g=2g{=}2 is often required to maintain performance at lower compute budget. Likewise, the advantage of g=2g{=}2 over g=1g{=}1 diminishes with difficulty, likely because harder tasks involve more reasoning errors, necessitating more frequent verifier intervention to stay on track.
Higher Number of Samples Tolerates Sparser Verification. In general, larger sample budgets allow for sparser verification, especially on easy to moderately difficult tasks. Higher budgets also amplify the performance gain of g=2g{=}2 over g=1g{=}1, as the generator can explore a wider search space, increasing the likelihood of finding a correct path despite less frequent verification.
@@ 5 Towards Adaptive Redefinition of the Verification Granularity
5.1 Adaptive Granularity Strategies
Our results (Figures 3, 5) show that the optimal $g$ in VG-Search depends on task difficulty, compute budget, and model capability. This motivates a more systematic approach to defining the search step, addressing RQ2. We formulate an optimization dual for selecting verification granularity: minimizing compute for a target accuracy vs. maximizing accuracy under a compute budget. For each, we propose a corresponding adaptive granularity strategy:
@@ 5.1 Adaptive Granularity Strategies
We compare against Beam Search, DVTS, and Best-of-N baselines on the MATH-500 test set. "val" indicates $g$ was tuned on a validation set MATH-250, while strategies under "oracle" use oracle $g$ selected on the test set. We set $\epsilon=0$ for CM-$g$. Table 1 shows that both AM-$g$ and CM-$g$ improve performance and efficiency. AM-$g$ consistently achieves higher accuracy, with gains up to 3.1% over Beam Search, while CM-$g$ provides substantial FLOPs savings, reducing compute by over 50% at $n=64$ and $n=256$ while maintaining or improving accuracy. For instance, CM-$g$ (val) at $n=128$ achieves 89.9% accuracy using just 11086 FLOPs—only 46% of the baseline budget. Although oracle-tuned variants perform slightly better, validation-tuned versions (CM-$g$ (val), AM-$g$ (val)) still outperform fixed-$g$ methods, demonstrating strong generalization and practicality. Under our cost model, $C_{\mathcal{G}}$ dominates, so most FLOPs savings come from pruning candidate paths during the Extend Step (Section 2.2). In summary, adapting verification granularity $g$ to task difficulty and compute budget offers a simple yet effective way, enabling more efficient and performant LLM reasoning.
5.2 Practicality and Convergence of Optimal Granularity Search
A key practical consideration is the overhead of finding the optimal granularity $g^{*}$. We analyzed its sample efficiency by selecting $g^{*}$ on validation subsets of increasing size and evaluating the performance on the MATH-500 test set. The results, shown in Table 2, demonstrate that the search converges rapidly. For both high ($n=256$) and medium ($n=64$) compute budgets, using just 5 validation samples achieves most of the possible accuracy gains over the baseline, confirming the minimal overhead of our adaptive approach.
This search constitutes a negligible, one-time cost that is amortized over all subsequent uses, ensuring our FLOPs comparisons are fair. The practical guideline is to reuse a pre-computed $g^{*}$ for a given model, task, and budget, and recompute it only when a core component, like the generator model, changes. For tasks with distinct subsets (e.g., varying difficulty levels), different $g^{*}$ values can be pre-computed and selected at inference time, making this a practical and effective alternative to the fixed-granularity convention.
Budget Strategy 0 samples 5 samples 100 samples 200 samples
$n=64$ CM-g 87.03% 88.57% 88.57% 88.57%
AM-g 87.03% 88.57% 89.19% 90.09%
$n=256$ CM-g 88.76% 90.37% 90.37% 90.43%
AM-g 88.76% 90.37% 90.37% 90.43%
Table 2: Test accuracy on MATH-500 vs. number of validation samples used to select $g^{*}$. Performance converges rapidly for both medium ($n=64$) and high ($n=256$) compute budgets.
@@ A.2 FLOPs Calculation using the Cost Model
where $L$ is the average solution length in generation steps, $T$ is the average number of tokens per generation step. To simplify compute tracking across experiments, we define a normalized compute proxy:
@@ A.3 Experiment Details for VG-Search
Baselines and Voting Methods. Our codebase is based on HuggingFace’s search-and-learn [4]. We adopt three mainstream TTS strategies: Best-of-N [5], Beam Search [32], and DVTS [4]. For all strategies we explore the number of samples of 4, 16, 64, 128, and 256. For Beam Search and DVTS, we use a fixed Branch-Out Factor $B_{2}=4$ following [32]. For all strategies, we use majority voting [37]. Temperature is set to 0.8 for all problems following HuggingFace’s setup.
@@
Hardware and Frameworks. We conduct our experiments on NVIDIA H100 [12] and A100 [11] GPUs, using CUDA 12.8 on Ubuntu 22.04. The vLLM library (v0.6.3) [18] is employed for model execution. The experiments for VG-Search on MATH-500 take from 1 to 5 hours, depending on the candidate number. To obtain accurate system-level runtime measurements, we use Nsight Systems [26] and employ the NVIDIA NVTX [13] extension to categorize execution time across different models. For the latency measurement in Figure 1(c), we use Qwen2.5-Math-1.5B as generator, while using RLHFlow/Llama3.1-8B-PRM-Deepseek-Data as the verifier.
A.4 Evaluation with an Ensemble Verifier
To broaden the scope of our evaluation, we conducted additional experiments using a multi-verifier ensemble. This approach can enhance verification quality by aggregating signals from diverse models [20]. For these experiments, we used Qwen-2.5-Math-1.5B as the generator. The verifier was an ensemble of two different Process Reward Models (PRMs): Skywork-1.5B [27] and RLHFlow’s Llama3.1-8B-PRM [39]. The final verification score for each reasoning step was computed by averaging the scores from these two models.
The results, presented in Table 3, align with and reinforce the main conclusion of our paper. We observe that the optimal verification granularity ($g$) is dynamic and highly dependent on the compute budget (i.e., the number of generations, $n$). While dense verification ($g=1$) is sometimes preferable at low sample counts, sparser verification ($g>1$) consistently achieves the highest accuracy on both the MATH-500 and AIME datasets as the compute budget increases. This confirms that our findings on adaptive verification granularity generalize even when using a stronger, ensembled verifier.
n MATH-500 (%) AIME (%)
g=1 g=2 g=3 g=4 g=1 g=2 g=3 g=4
4 76.0 73.2 71.6 69.6 12.22 8.89 11.11 13.33
16 80.0 77.4 76.4 78.4 13.33 13.33 16.67 14.44
64 79.4 80.4 80.8 81.0 16.67 16.67 16.67 18.89
128 80.8 82.2 82.2 80.8 15.56 24.44 23.33 18.89
256 80.2 80.0 80.2 81.2 23.33 23.33 23.33 23.33
Table 3: Accuracy of VG-Search with an ensemble verifier across MATH-500 and AIME datasets. For each number of generations ($n$), the highest accuracy in each row is bolded, indicating the performance achieved at the optimal granularity $g$.
A.5 Applicability to Tasks with Outcome-Level Feedback
VG-Search is applicable to tasks with outcome-level feedback, provided a PRM can be trained on intermediate generation steps. In such scenarios, the core search algorithm remains unchanged. However, for open-ended generation tasks like code synthesis where outputs are diverse, standard majority voting is not feasible for final answer selection. A practical alternative is to select the single candidate trajectory with the highest cumulative or final PRM score.
To demonstrate this, we evaluated VG-Search on the HumanEval benchmark using LLaMA-3.2-1B as the generator and Skywork-o1-1.5B-PRM as the verifier. We used newline characters to delimit generation steps. The results, reported as pass@1 scores in Table 4, show that sparser verification ($g>1$) can outperform dense verification ($g=1$), particularly as the compute budget ($n$) increases. This confirms that our central conclusion—that the optimal verification granularity is dynamic—generalizes to open-ended generation tasks.
n g=1 g=2 g=3 g=4
4 0.250 0.274 0.238 0.220
16 0.335 0.305 0.268 0.311
64 0.256 0.335 0.274 0.305
Table 4: Pass@1 scores on HumanEval using VG-Search, selecting the highest-scoring candidate. The best performance for each compute budget ($n$) is bolded.
A.6 Standard Deviation with Multiple Trials
To evaluate the robustness of our results, we conducted experiments over multiple trials using different random seeds. This analysis provides a more rigorous assessment of variance and confirms the stability of our conclusions. The following results were generated using the Qwen-Math-7B generator and the Skywork-o1-1.5B PRM on the MATH-500 dataset, averaged over three distinct runs.
Table 5 shows the mean accuracy and standard deviation for this experiment. The results from multiple trials are consistent with our original findings, demonstrating the stability of the observed trends. The low standard deviations across all configurations reinforce this conclusion. While dense verification ($g=1$) performs well at smaller compute budgets, sparser verification ($g>1$) consistently achieves higher accuracy as the number of samples increases ($n\geq 64$).
n g=1 g=2 g=3 g=4
4 84.13 $\pm$ 0.50 82.27 $\pm$ 0.24 81.73 $\pm$ 0.39 80.40 $\pm$ 0.43
16 86.80 $\pm$ 0.28 86.13 $\pm$ 0.22 86.13 $\pm$ 0.25 86.40 $\pm$ 0.16
64 86.20 $\pm$ 0.33 87.27 $\pm$ 0.15 88.27 $\pm$ 0.34 87.87 $\pm$ 0.19
128 86.67 $\pm$ 0.19 87.60 $\pm$ 0.57 88.27 $\pm$ 0.21 88.47 $\pm$ 0.32
256 86.67 $\pm$ 0.19 87.93 $\pm$ 0.34 89.13 $\pm$ 0.27 88.73 $\pm$ 0.19
Table 5: Mean accuracy $\pm$ standard deviation over 3 runs with different seeds on MATH-500. The highest mean accuracy for each sample count ($n$) is bolded.
A.7 Limitations
@@
A.8 Case Study of VG-Search